OpenStudy (anonymous):

flying to warsaw with a tailwind a plane averaged 231 km/h. on the return trip the plan only averaged 173 km/h while flying back into the same wind. find the speed of the plane in still air and the speed of the wind.

OpenStudy (amistre64):

speed = distance/time. since the distance stays the same, we can solve for the rest of it....somehow :)

OpenStudy (anonymous):

o okay

OpenStudy (amistre64):

As + Ws = 231 As - Ws = 173 As = 231 - Ws and As = 173 + Ws in other words: 231 - Ws = 173 + Ws Solve for Ws....

OpenStudy (amistre64):

231 - 173 = Ws + Ws 58 = 2Ws 58/2 = Ws 29 = Ws The Wind speed should be 29 km/hour

OpenStudy (amistre64):

As + 29 = 231 As = 231 - 29 Air speed = 202 km/hr best double check me work tho :)

OpenStudy (anonymous):

thank uu! (: